Two particle comes towards each other with a speed 0.8c with respect to the laboratory their relative speed is
1
Expert's answer
2019-07-04T11:10:15-0400

in this case, the relative speed

"u=\frac{2v}{1+(\frac{v}{c})^{2}}=\frac{1.6c}{1+0.64} = 0.98c"